# Time Distributed Deep Learning models for Purely Exogenous Forecasting. Application to Water Table Depth Prediction using Weather Image Time Series

### v1

@@ Abstract
Groundwater resources are one of the most relevant elements in the water cycle, therefore developing models able to accurately predict them is a pivotal task in the sustainable resources management framework. Deep Learning (DL) models have been revealed very effective in hydrology, especially by feeding directly spatially distributed data (e.g. raster data). In many regions, hydrological measurements are difficult to obtain regularly or periodically in time, and in some cases, last available data are not up to date. Reversely, weather data, which have significant impacts on water resources, are usually more available and with higher quality. More specifically, we have proposed two different DL models to predict the water table depth in the Grana-Maira catchment (Piemonte, IT) using only exogenous weather image time series. To deal with the image time series, both the model are made of a first Time Distributed Convolutional Neural Network (TDC) which encodes the image available at each time step into a vectorial hidden representation. The first model, TDC-LSTM uses then a Sequential Module based on an LSTM layer to learn temporal relations and output the predictions. The second model, TDC-UnPWaveNet uses instead a new version of the WaveNet architecture, adapted here to output a sequence shorter and completely shifted in the future with respect to the input one. To this aim, and to deal with the different sequence lengths in the UnPWaveNet, we have designed a new Channel Distributed layer, that acts like a Time Distributed one but on the channel dimension, i.e. applying the same set of operations to each channel of the input. TDC-LSTM and TDC-UnPWaveNet have shown both remarkable results. However, the two models have focused on different learnable information: TDC-LSTM has focused more on lowering the bias, while the TDC-UnPWaveNet has focused more on the temporal dynamics maximising correlation and KGE.
@@ 1 Introduction
Water resources management is a key element in the sustainable development framework, even more in the case of climate change [1]. Freshwater is essential for sanitation and hygiene standards but also for granting food availability and economic stability [2]. Groundwater resources are the second most relevant component of the water cycle, accounting for roughly 30% of freshwater resources all over the world, just after glacier and ice caps which account for 68.7% [3]. Furthermore, groundwater resources prove to be a more stable source of freshwater with respect to others like rivers and lakes, whose water storage varies promptly with the current weather conditions [4]. In fact, In Europe, 65% of drinking water and 25% of irrigation comes from groundwater resources [5] and these percentages are expected to increase given the expected effects of climate change [1, 6, 2, 7].
Water management policies are a pivotal tool to pursue and achieve sustainable development [8, 9] and fulfil the needs of water both in terms of drinkable water for humans and irrigation of crops. Accounting for the water resources of a region is of utmost relevance to plan water policies, and for this, it is essential to develop models which can quantify groundwater resources. The phenomena concerning groundwater are extremely complex and very related to the physical properties of the context [10, 11], not by chance physical models need a lot of information of the region of concern and are very scale-dependent [12, 13]. In the last years, Deep Learning (DL) techniques have been applied to groundwater modeling, and they have proved to be very efficient in obtaining remarkable results without needing all the physical knowledge of the region under study [14, 15, 16, 17, 18, 19, 20, 21]. In particular, in addition to autoregressive terms providing information on the past states of the variable to predict (target variable), most of the DL studies use exogenous weather variables (e.g. precipitation, temperature, etc.) as the major input motivated by domain knowledge and data availability. In fact, other types of data, like the anthropogenic pressure on water resources (e.g. water abstractions, irrigations), are used only by a minority of studies because of the scarce availability in many catchments, and in general at the large scale [22, 23, 24].
@@
To quantify groundwater resources, groundwater level (GWL) is usually adopted as the target variable [20]. GWL represents the distance between the sea level and the higher surface of the groundwater body, which often is defined as the water table, i.e. the surface of the phreatic aquifer. Another possibility of quantification is to measure the water table depth, i.e. the distance between the ground surface and the water table. In this way, a decrease in the water table depth means an increase in groundwater resources, reversely an increase of the water table depth means a decrease of the water stored in the phreatic aquifer. The more widespread data type adopted to predict the water table depth is tabular data (i.e one-dimensional time series) for both the input and the output, as in [25, 19, 21]. However, [26, 27] revealed the usefulness of using spatially distributed input data (e.g. raster data) as input data for hydrological data-driven modeling, letting the DL model find the most useful relations among all the input variables spread over the region of interest (ROI). This could be done using as input, a time series of weather raster images (i.e. a multidimensional time series), in which an image for each considered weather variable is retrieved [^1] In case of more weather variables one can speak of a multivariate and multidimensional time series. at each time step $t$. The values of a pixel in an image is the vector of values associated with the corresponding weather variables for the area covered by that pixel.
@@
Spatially distributed data (i.e. images) could be efficiently and effectively handled by 2D Convolutional Neural Networks (CNN), which have been extensively proficiently adopted in DL models to deal with images and learn spatial relations [28, 29, 30, 31, 32]. CNN has been explored also for learning temporal relations, i.e. sequential data; for example, 1D CNN has been widely adopted [33] also in water resources studies [25, 34, 20]. However, Recurrent Neural Networks (RNN) have proved to be a very hard-to-beat competitor for sequential data, especially Long-Short-Term Memory (LSTM) architecture thanks to its ability to learn both short and long-term dependencies being resilient to noise [35, 36]. In the last year, many researchers have tried to build new and more sophisticated CNN-based architectures to compete with LSTM on sequential data [37, 38, 39, 40, 41, 42, 43, 44]. These strengths are justified by the fact that CNNs are less computational intensive and highly parallelizable, meaning that their training is less energy consuming. Furthermore, with the use of dilated convolution (also named ”à trous convolution”) [45], it has been possible to implement CNN based model with the ability to capture very long-term relations with a lower number of parameters. Different architectures have been proposed adopting the dilated convolution, and they have shown very competing performances in comparison to RNN [37, 39]. One example could be the WaveNet model, developed by Google [46] for audio generation in a many-to-many framework [^2] In computer science there are different framework concerning sequential data modeling: many-to-one (or seq2one) for models which take as input a sequence and output a scalar, many-yo-many (or seq2seq) for models which take as input a sequence and output a sequence.. WaveNet uses dilated convolution to learn long dependencies, and it adopts causal padding to constrain each element of the output sequence to depend on only past input observations (see Section 2 for more details). This architecture has been widely used also for other task than audio generation producing remarkable results [40, 47, 44].
@@
To deal with image time series many studies combined 2D CNN and a sequential model (e.g. LSTM or 1D CNN) to get the most from them. More in detail, 2D CNN has been used in a Time Distributed (TD) way, i.e. applying the same 2D CNN to the image available at each time step and extracting spatial relations from it. Then, the output of the TD 2D CNN is fed to a sequential model which focuses on the temporal relation of the data - sometimes in the literature, these models are referred as hybrid models (e.g. CNN-LSTM) [48, 49, 50, 51, 52, 53, 27, 26]. As already stated, an image time series of many weather variables could be seen as a multidimensional (time, longitude, and latitude dimensions) and multivariate (a single pixel is a vector of variables) time series. However, from a more computer science viewpoint, this data flow could be seen as a video, in which per each frame (i.e. time-step) many channels (i.e. variables or features) are present. In fact, many studies adopt these hybrid models for different video tasks [54, 55, 56]. To be readable by the two communities, in the following, we will use the term channels interchangeably with variables, and image time series with video.
@@
Regarding the application, the present research focuses on the Grana-Maira catchment in Piemonte (Italy). Our goal, a many-to-one task, is to forecast the weekly water table depth measured by three sensors in the catchment area (i.e. our ROI) feeding as input, exogenous weather information of the previous two years in the form of image time series over the ROI. To do this, we developed two different DL models. Inspired by literature, both models are made by two modules. The first module, named Time Distributed CNN (TDC in short) is the same for both models, and it is responsible for handling the images available at each time step of the image time series and learning spatial relations. What distinguishes the two models is the second module (Sequential Module hereafter), which is responsible for learning temporal relations. The first model, named TDC-LSTM, has a Sequential Module based on an LSTM layer. Differently, the second model, named TDC-UnPWaveNet, uses a new version of the WaveNet, here proposed to be usable for the many-to-one case, and in general, for tasks in which the output sequence has a lower length and it is completely shifted in the future with respect to the input sequence. This adaptation has required a restructuring of the original WaveNet model and the development of a new Channel Distributed (CD) layer for dealing with objects of different time lengths between the hidden layers of the architecture.
@@ 1.1 Related Works
DL techniques have already been applied proficiently to groundwater resource forecasting. In [25] authors made a comparison of different DL architectures, namely NARX [57], 1D CNN and LSTM to predict GWLs on 17 sensors in the Upper Rhine Graben (URG) region. NARX is a neural network architecture specifically designed to model autoregressive terms (i.e. past values of the target) and exogenous data in a nonlinear fashion. They trained local models for each sensor using Bayesian optimization and fed as input, weather variables measured by nearby sensors (i.e. in a tabular structure), furthermore, for some of the sensors, also an autoregressive term. They found CNN faster in training and inference than other methods. LSTM was revealed to be the worst performing, while NARX performed the best, but this could be thanks to the intrinsic use of an autoregressive component, not explicitly provided to CNN and LSTM. However, in many other works, LSTM performed better than other methods like ARIMA, ANN[58], Random Forest [18]. In [59] authors made a comparison over many models and they found LSTM and NARX as the best models without a clear winner between the two. Notwithstanding, LSTM seems to be more established and widespread in the DL community and also in hydrological applications, which more frequently adopt LSTM as the reference model for data-driven modeling [60, 61, 62, 63].
@@
Most of the studies on groundwater modeling deals with tabular data to model their target. This means retrieving input and output data directly from the measurement sensor network over the region of interest (i.e. geospatial data in more statistical terms). In this way, each observation in the dataset is indexed by the time of acquisition, and longitude and latitude of the sensor by which it is measured. Spatially distributed data instead are represented in a grid format (i.e. raster) in which each portion of the area under study is represented by a square of the grid (or a pixel if one looks at the raster as an image). This type of data represents the variables of interest all over the region under study, and not only on the coordinate in which the sensors are located. Spatially distributed data could be created either by spatial interpolation of tabular data or by using other sensors like satellites (e.g. GRACE data [64]). In both cases, in the DL framework, using spatially distributed data as input could facilitate the model in understanding spatial relations and extracting the most relevant information without making a priori assumptions on the form of these spatial relations.
@@
Some works in hydrology retrieved spatially distributed data, however, many of them squeezed the spatial dimension before feeding the data to the model, for example by averaging the value of the variable over the watershed of interest [65, 21]. Instead, in [27, 26] authors used spatially distributed weather data (i.e. raster images), as direct input to model streamflow in Canada and spring discharge in karst catchments respectively. In general, the use of spatially distributed data could be helpful for two reasons. The first is that thanks to new open raster datasets like ERA5-land [66], it is possible to retrieve water and energy-related variables also in regions which have not an extensive measurement network of sensors. The second concerns the possibility of building complex models that autonomously learn the most relevant spatial relations between the weather input variables and the target. In more detail, [27] propose a neural network made by 2 modules to jointly predict streamflow for 226 stream gauge stations. The first module is a TD CNN made of 5 convolutional layers and 2 max pooling layers. The input images at each time step are squeezed in the spatial dimension obtaining a vector of dimension 32, then the input video is converted into a multivariate (hidden) time series with 32 features. The second module is a one-layer LSTM with 80 units, and the last is a fully connected layer which outputs the predictions for the 226 stream gauges. In [25], authors made a comparison between the tabular and spatially distributed approaches in three different areas. For the spatially distributed data they developed still a 2 modules model whose first module is a TD CNN, but they used a 1D CNN layer as the second module. They argued that by using spatially distributed data, one can overcome the difficulties in areas with a low density of measurement sensor. Furthermore, for studies focused on more catchments, and the aid of sensitivity analysis, it is possible to extract a naive localization of the catchment by looking at the most sensitive pixel in the original raster images - a practice belonging to the perturbation methods in the eXplainable AI (XAI) research field [27, 67].
@@
To the best of our knowledge, no other studies have tried to feed spatially distributed data directly into a DL model to predict the water table depth, especially in our ROI and making a direct comparison of hybrid recurrent vs convolutional methods. Thus, our contributions are the following: a) development of DL models for water table depth predictions in the Grana-Maira catchment in Piemonte (IT) b) development and application of DL hydrological models which directly use spatially distributed data c) adaptation of WaveNet to the many-to-one case through the UnPWaveNet, a new competitor to recurrent architectures for sequential data.

### v2

# Time Distributed Deep Learning Models for Purely Exogenous Forecasting: Application to Water Table Depth Predictions using Weather Image Time Series

@@ Abstract
Deep Learning (DL) models have revealed to be very effective in hydrology, especially in handling spatially distributed data (e.g. raster data). We have proposed two different DL models to predict the water table depth in the Grana-Maira catchment (Piedmont, IT) using only exogenous weather image time series. Both the models are made of a first Time Distributed Convolutional Neural Network (TDC) which encodes the images into hidden vectors. The first model, TDC-LSTM uses then a Sequential Module based on an LSTM layer to learn temporal relations and output the predictions. The second model, TDC-UnPWaveNet uses instead a new version of the WaveNet architecture, adapted for handling output of different length and completely shifted in the future to the input. Both models have shown remarkable results focusing on different learnable information: TDC-LSTM has focused more on bias while the TDC-UnPWaveNet more on the temporal dynamics maximizing correlation $\rho$, achieving mean BIAS (and standard deviation) -0.18(0.05), -0.25(0.19) and $\rho$ 0.93(0.03), 0.96(0.01) respectively over all the sensors.
@@ 1 Introduction
Water management is a key element in the framework of sustainable development, even more so in the context of climate change [1]. Freshwater is essential for sanitation and hygiene standards, but also for food availability and economic stability [2]. Groundwater resources are the second most important component of the hydrological cycle, accounting for about 30% of the world’s freshwater resources, just behind glaciers and ice caps, which account for 68.7% [3]. Moreover, groundwater resources prove to be a more stable source of freshwater compared to rivers and lakes, whose water storage varies promptly with the current weather conditions [4]. Indeed, in Europe, 65% of drinking water and 25% of irrigation comes from groundwater resources [5]; these percentages are expected to increase given the expected effects of climate change [1, 6, 2, 7].
Water management policies are a key tool to pursue and achieve sustainable development [8, 9] and to meet water needs both in terms of drinking water for humans and irrigation for crops. Thus, accounting for water resources is paramount for planning water policies, and it is essential to develop models that can quantify groundwater resources.
The traditional process-based approaches consist of developing hydrological models mapping explicitly the cause-effect relationship among the variables under study. Different solutions have been proposed ranging from conceptual to physically-based and distributed models [10, 11]. However, calibrating process-based model parameters requires many geophysical data of the aquifer under study and a deep knowledge of the hydrogeophysical processes involved [12, 13]. This turns into costly and time-consuming field measurements, that in some developing rural areas are even unfeasible [14, 15, 16, 17]. Furthermore, even if computing resources have grown substantially, they are still stretched to their limits by simulation of complex spatio-temporal hydrological phenomena [18]. Indeed, execution times of simulations remain an issue hindering the use of accurate process-based models for nowcasting applications. Nevertheless, process-based models remain a cornerstone for developing explainable models that obey physical laws, and in cases in which the possibility of explaining how the model produces a specific output (i.e. having a white box model) is more relevant than having accurate and fast predictions. Authors in [19] proposed a new mathematical method to improve the modeling of surface–groundwater interactions while still maintaining a formal description of the model functioning. In more detail, they adopted spectral analysis to improve the accuracy of the Fourier fitting process to identify the significant surface signals. The approach was tested on synthetic data showing improved accuracy in approximating surface and boundary fluxes in the hydrological model.
In the last years, different, and fully empirical, approaches have emerged leveraging the growing availability of hydrological measurement data. In more detail, machine learning techniques have been applied to groundwater modeling and have proven to be very efficient in achieving remarkable results without any prior domain knowledge, neither in terms of physics laws nor in terms of aquifer geophysical details [20, 21, 22]. Deep Learning (DL) models, based on neural networks, have reached even better performances thanks to their ability to learn very complex and nonlinear relations between input and output [23, 24, 17, 25, 10, 11]. Indeed, neural networks have been demonstrated to be universal approximators [26], and thus they can approximate any continuous function with a sufficient number of model parameters and an adequate amount of data. A critical aspect of deep learning is the computational cost of training neural networks. However, highly optimized libraries and Graphical Processing Units (GPU) that speed up the execution of parallelized algorithms help in reducing training times. Furthermore, it is worth considering that once a model is trained, predictions could be obtained rapidly, promoting the adoption of DL for nowcasting [27].
Looking more at the implementation viewpoint, DL models for groundwater application have been developed by using mainly exogenous weather data as input, and in some cases, some autoregressive terms (past target data) have been added as input to improve performances. Other types of data, such as anthropogenic pressure on water resources (e.g. water abstraction, irrigation), are only used by a minority of studies due to their scarce availability in many catchments and at large scales [28, 29, 30].
@@
To quantify groundwater resources, groundwater level (GWL) is frequently adopted as the target variable [21]. GWL represents the distance between the reference datum and the water table, which in turn is defined as the higher surface of the groundwater body. Another possibility of quantification is to measure the water table depth, i.e. the distance between the ground surface and the water table. In this way, a decrease in the depth of the water table means an increase in groundwater resources; conversely, an increase in the depth of the water table means a decrease in the amount of water stored in the phreatic aquifer. The more widely used data type for groundwater depth prediction is tabular data (i.e. one-dimensional time series) for both input and output, as in [31, 25, 22]. However, the authors of [32, 33] demonstrated the usefulness of using spatially distributed input data (e.g. raster data) as input data for hydrological data-driven modeling, allowing the DL model to find the most useful relationships among all input variables distributed over the region of interest (ROI). This could be done by taking as input a time series of weather raster images (i.e. a multidimensional time series), where for each weather variable an image is retrieved at each time step $t$. The values of a pixel in an image are the vector of values associated with the corresponding weather variables for the area covered by this pixel.
@@
Spatially distributed data (i.e. images) could be efficiently and effectively handled by 2D Convolutional Neural Networks (CNNs), which have been extensively and proficiently adopted in DL models to deal with images and learn spatial relationships [34, 35, 36, 37, 38]. CNNs have been explored also for learning temporal relationships, i.e. sequential data; for example, 1D CNNs have been widely adopted [39] also in water resources studies [31, 40, 21]. However, Recurrent Neural Networks (RNNs) and Long-Short-Term Memory (LSTM) have proven to be very hard to beat competitors for sequential data, thanks to their ability to learn both short and long term dependencies and their resilience to noise [41, 42]. In recent years, many researchers have attempted to build new and more sophisticated CNN-based architectures to compete with LSTM on sequential data [43, 44, 45, 46, 47, 48, 49, 50]. These strengths are justified by the fact that CNNs are less computationally intensive and highly parallelizable, meaning that their training is less energy-consuming. Furthermore, with the use of dilated convolution (also called ”hole convolution”) [51], it has been possible to implement CNN-based models with the ability to capture very long-term relationships with a smaller number of parameters. Different architectures have been proposed adopting the dilated convolution, and they have shown very competing performances in comparison to RNN [43, 45]. One example could be the WaveNet model, developed by Google [52] for audio generation in a many-to-many framework [^1] In machine learning literature there are different frameworks for sequential data modeling: many-to-one (or seq2one) for models which take as input a sequence and output a scalar, many-to-many (or seq2seq) for models which take as input a sequence and output a sequence.. WaveNet uses dilated convolution to learn long dependencies and employs causal padding to constrain each element of the output sequence to depend only on past input observations (see Section 2 for more details). This architecture has been widely used for tasks other than audio generation, with remarkable results [46, 53, 50].
@@
To deal with image time series many studies combined 2D CNN and a sequential model (e.g. LSTM or 1D CNN) to get the most from them. More specifically, 2D CNNs have been used in a time-distributed (TD) manner, i.e. the same 2D CNN is applied to the image available at each time step, and spatial relationships are extracted from it. The output of the TD 2D CNN is then fed into a sequential model that focuses on the temporal relationships of the data - sometimes these models are referred to in the literature as hybrid models (e.g. CNN-LSTM) [54, 55, 56, 57, 58, 59, 33, 32]. As mentioned above, from a statistical viewpoint, an image time series of many weather variables can be seen as a multidimensional (in the dimensions of time, longitude and latitude) and multivariate (more than one input variable) time series. Instead, from a Computer Science point of view, this data flow could be seen as a video in which each frame (i.e. time step) contains many channels (i.e. variables or features). Indeed, many studies adopt these hybrid models for different video tasks [60, 61, 62]. To be readable by the two communities, in the following, we will use the term channels interchangeably with variables, and image time series with video.
@@
Regarding the application, the present research focuses on the Grana-Maira catchment in Piedmont (Italy). Our objective, a many-to-one task, is to predict the weekly water level depth measured by three sensors in the catchment area (i.e. our ROI). The input consists of exogenous weather information from the last two years and takes the form of image time series over the ROI. To this aim and inspired by literature, we developed and compared two different composite (i.e. made of sub-modules) DL models. The first sub-module, named Time Distributed CNN (TDC) is the same for both, it is responsible for handling the images available at each time step of the time series while learning spatial relations. What distinguishes the two models is the second module (hereafter Sequential Module), which is responsible for learning temporal relations. The first model, named TDC-LSTM, has a Sequential Module based on an LSTM layer. Differently, the second model, named TDC-UnPWaveNet, uses a new version of the WaveNet, which is proposed here to be usable for the many-to-one case and, in general, for tasks in which the output sequence has a lower length and it is completely shifted in the future to the input sequence. This adaptation required a restructuring of the original WaveNet model and the development of a new Channel Distributed (CD) layer to handle objects of different time lengths between the hidden layers of the architecture.
@@ 1.1 Related Works
DL techniques have already been applied proficiently to groundwater resource forecasting. In [31] authors made a comparison of different DL architectures, namely NARX [63], 1D CNN, and LSTM to predict GWLs on 17 sensors in the Upper Rhine Graben (URG) region. NARX is a neural network architecture specifically designed to model autoregressive terms (i.e. the past values of the target) and exogenous data in a nonlinear fashion. They trained local models for each sensor using Bayesian optimization and fed as input weather variables measured by nearby sensors (in other words inputs had a tabular structure); furthermore, for some of the sensors, there was also an autoregressive term. The authors found that CNN was faster in training and inference than other methods. LSTM was revealed to be the worst performing, while NARX performed the best, but this conclusion could be due to the intrinsic use of an autoregressive component by the NARX, not explicitly provided to CNN and LSTM. Nevertheless, in many other works, LSTM performed better than other methods like ARIMA, ANN [64], and Random Forest [17]. In [65], the authors made a comparison across many models and found that LSTM and NARX were the best models, with no clear winner between the two. However, LSTM seems to be more established and widespread in the DL community and also in hydrological applications, which more often adopt LSTM as a reference model for data-driven modelling [66, 67, 68, 69].
@@
Most of the studies on groundwater modeling deal with tabular data to model their target. This means retrieving input and output data directly from the measurement sensor network over the region of interest (i.e. geospatial data in more statistical terms). In this way, each observation in the dataset is indexed by the time of acquisition, and longitude and latitude of the sensor by which it is measured. Spatially distributed data instead are represented in a grid format (i.e. raster) in which each portion of the area under study is represented by a square of the grid (or a pixel if one looks at the raster as an image). This type of data represents the variables of interest all over the region under study, and not only on the coordinates in which the sensors are located. Spatially distributed data could be created either by spatial interpolation of tabular data or by using other sensors like satellites (e.g. GRACE data [70]). In both cases, in the DL framework, using spatially distributed data as input could facilitate the model in understanding spatial relationships and extracting the most relevant information without making a priori assumptions on the form of these spatial relations.
@@
Some works in hydrology retrieved spatially distributed data, however, many of these have reduced the spatial dimension before feeding the data into the model, for example by averaging the value of the variable over the watershed of interest [71, 22]. Instead, the authors in [33, 32] used spatially distributed weather data (i.e. raster images) as direct input to model streamflow in Canada and spring discharge in karst catchments, respectively. In general, the use of spatially distributed data could be helpful for two reasons. The first is that, thanks to new open raster datasets such as ERA5-land [72], it is possible to retrieve water and energy-related variables even in regions that do not have an extensive sensor network. The second is the ability to build complex models that autonomously learn the most relevant spatial relationships between the weather input variables and the target. In more detail, in [33] a neural network made by 2 modules is proposed to jointly predict streamflow for 226 stream gauge stations. The first module is a TD CNN made of 5 convolutional layers and 2 max pooling layers. The input images at each time step are squeezed in the spatial dimension to obtain a vector of dimension 32, then the input video is transformed into a multivariate (hidden) time series with 32 features. The second module is a one-layer LSTM with 80 units, and the last is a fully connected layer that outputs the predictions for the 226 stream gauges. In [31], the authors made a comparison between the tabular and spatially distributed approaches in three different areas. For the spatially distributed data, they still developed a 2-module model, where the first module is a TD CNN, and the second module is a 1D CNN layer. They argued that by using spatially distributed data, one can overcome the difficulties in areas with a low density of sensors. Furthermore, for studies focused on more catchments, and with the aid of sensitivity analysis, it is possible to extract a naive localization of the catchment by looking at the most sensitive pixels in the original raster images - a practice belonging to the perturbation methods in the eXplainable AI (XAI) research field [33, 73].
@@
To the best of our knowledge, no other studies have attempted to feed spatially distributed data directly into a DL model to predict the water table depth, especially in our ROI, and to make a direct comparison between hybrid recurrent and convolutional methods. Thus, our contributions are the following:
  1. 1.
    Development of DL models for water table depth predictions in the Grana-Maira catchment in Piedmont (IT)
  2. 2.
    Development and application of DL hydrological models that directly use spatially distributed data
  3. 3.
    Adaptation of WaveNet to the many-to-one case through UnPWaveNet, a new competitor to recurrent architectures for sequential data.